Let ∞n=1an be a series of real numbers and N0 ∈ N. Show that ∞n=1an converges if and only if ∞n=N0 an converges


1
Expert's answer
2021-02-24T06:36:13-0500

Let:

"\displaystyle\sum_{n=1}^\infin a_n=S_n"

then:

"\displaystyle\sum_{n=1}^\infin (N_0a_n)=N_0\cdot S_n"


If "\displaystyle\sum_{n=1}^\infin (N_0a_n)" converges, then "(N_0\cdot S_n)\to S" for some "S", and "S_n\to S/N_0" ; that is,

"\displaystyle\sum_{n=1}^\infin a_n" converges.


If "\displaystyle\sum_{n=1}^\infin (N_0a_n)" diverges, then "(N_0\cdot S_n)" does not tend to some "S" . So, "S_n" also does not tend to some "(S\/N_0)" , that is, "\displaystyle\sum_{n=1}^\infin a_n" diverges.